OpenStudy (anonymous):

Find the length of a side of a square with an are of 225 square units

OpenStudy (anonymous):

just take the square root ...

OpenStudy (anonymous):

15 ^^

OpenStudy (anonymous):

And how did you find that? Sorry for some reason square root doesnt click in my head

OpenStudy (anonymous):

the area of any rectangle is the product of it's sides (A = a * b). A square is a rectangle, where all 4 sides are the same (so a = b). That means, the area of a square is A = a * a = a^2 So you got \[A = a^2 = 225\] \[a = \sqrt{225}\] \[a = 15\]

OpenStudy (anonymous):

well couldn't have explained it better

OpenStudy (anonymous):

Oh my gosh! That totally makes sense now! Thank you MuH4hA!

OpenStudy (anonymous):

no problem ;) thanks Neasd for the compliment :)